a farmhouse has 15 animals. Some are goats and some are ducks. altogether there are 48 legs. how many of each animal are there?
Answer by checkley77(12569)

D+G=15 OR D=15-G
2D+4G=48
2(15-G)+4G=48
30-2G+4G=48
2G=48-30
2G=18
G=18/2
G=9 NUMBER OF GOATS.
15-9=6 DUCKS.
PROOF:
9*4+6*2=48
36+12=48
48=48